# JEPA-Reasoner: Decoupling Latent Reasoning from Token Generation

### v2

@@ 5.2.3 Model Configurations and Training Methods
We denote the vanilla Transformer baseline as $T$, the COCONUT-style coupled latent space reasoning model as $C$, and the decoupled model as $R$ (both JEPA-Reasoner and Talker are included). Table 3 shows detailed model configurations used in this section:
@@
We first pretrain Transformer models on CFG data with cross-entropy loss, then pick the checkpoint with optimal accuracy on CFG completion task. This unified checkpoint serves as the initialization for all subsequent models. [^2] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count. We then post-train the Transformer to obtain baseline $T$. COCONUT models are trained to predict a sequence of four hidden states followed by four tokens, with loss computed exclusively on the target tokens. The JEPA-Reasoner and Dual Talker model $R$ are trained following the methodology in Section 4. All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, effective batch size 128, context length 1024) and all models are trained until their loss stabilizes. The best-performing checkpoint is selected as the representative for each architecture.
@@ 5.2.4 Results and Conclusions
Our robustness evaluation demonstrates the advantages of the decoupled architecture. In the token-level error experiment, JEPA-Reasoner demonstrated less performance degradation when facing input noise during multi-step CFG completion tasks (Figure 2). JEPA-Reasoner also exhibits higher performance across different magnitudes of Gaussian noise in the latent space error experiment (Table 4), providing more empirical evidence for its robustness advantage.
@@
Table 4: Performance of JEPA-Reasoner $R$ and COCONUT $C$ under different latent noise levels.
@@
These results demonstrate that JEPA-Reasoner has the potential to address the limitations of existing paradigms in Section 2. By operating in a normalized latent space and offloading token generation to the Talker module, subsequent reasoning outputs do not condition on previous decisions, thereby mitigating error accumulation in the autoregressive process and enabling more robust sequential generation under noisy conditions.
@@ 6 Real World Evaluation
We trained a 694M JEPA-Reasoner and a 198M Mono-Talker on natural language using the training steps stated in Section 4. We analyze the performance in two ways:
@@
    Performance difference between pretrained model (token level prediction) and JEPA-Reasoner (latent level prediction), paired with its Mono-Talker model.
@@
    Performance difference between JEPA-Reasoner (paired with Mono-Talker) and other latent or non-latent reasoning models.
6.1 Performance Gain After Latent Reasoning Adaptation
We evaluated both the base Transformer model and JEPA-Reasoner on GSM8K (gsm8k). Table 5 shows the benchmark score under both 5-shot and 8-shot settings.
Table 5: GSM8K benchmarks of base Transformer and JEPA-Reasoner
@@ 6.1 Performance Gain After Latent Reasoning Adaptation
Performance Gain.
Comparing to the base Transformer, JEPA-Reasoner improves performance by 79.2% (5-shot) and 149.5% (8-shot), respectively, demonstrating that adapting to a decoupled latent space reasoning paradigm significantly increases the accuracy of expressing internally learned knowledge with explicitly stated token sequences. Considering that it takes 300k steps for pretraining but only 13k steps for SST, while using the same dataset, it shows great evidence that most of the performance gain came from the adaptation to latent reasoning, rather than SST steps. Also, the scaled cosine similarity loss used in SST encourages smooth latent representation transition rather than logical correctness, which further support that minimal new knowledge was learned during SST.
@@
Crucially, while the Base Transformer’s performance stagnates between 5-shot and 8-shot settings (improving only by 0.1%), the JEPA-Reasoner utilizes the additional examples to nearly double its accuracy. This scaling behavior indicates that the decoupled latent space reasoning architecture effectively overcomes the reasoning plateau, which was widely observed in small-scale token-based models, allowing for more robust logical deduction that is less constrained by surface-level token statistics.
6.2 Performance Comparison with Other Reasoning Models
We compare JEPA-Reasoner against standard Transformer models and other reasoning models. Table 6 shows the GSM8K benchmark performance of our JEPA-Reasoner model (paired with Talker) compared to other models of similar size or performance.
Table 6: GSM8K benchmark comparison between JEPA-Reason and other reasoning models. Note that we could not find 8-shot GSM8K benchmark results for COCONUT models. Since COCONUT models sacrifice performance (compared to CoT) as a trade-off for efficiency (coconut), we did not include them in the performance comparison.
@@ 6.2 Performance Comparison with Other Reasoning Models
Ours JEPA-Reasoner 0.9B 51.9
@@
As illustrated in Table 6, JEPA-Reasoner demonstrates superior performance on the GSM8K benchmark compared to both standard transformer baselines and dedicated CoT models. With a parameter count of only 0.9B, our model achieves an 8-shot score of 48.2%, outperforming Llama 3.2 (1B) CoT baseline by 3.8 percentage points. Most notably, JEPA-Reasoner significantly surpasses larger models that rely on standard reasoning paradigms. Despite being approximately $4\times$ smaller than Gemma 3 (4B), our model exhibits a performance gain of nearly 10%. While the Recurrent Depth model (Huginn-0125) offers a strong baseline at 42.1%, it requires nearly four times the parameter count to achieve results that are still 6.1% lower than JEPA-Reasoner. Consequently, these results serve as strong empirical evidence of JEPA-Reasoner’s capability to handle complex natural language reasoning tasks, effectively applying mathematical logic to solve problems.
@@ 7 Theoretical Advantage of Decoupled Architecture
The robustness of our decoupled architecture stems from its ability to decouple the high-level reasoning process from low-level token generation. We can formalize this advantage by analyzing the probabilistic assumptions and information flow within the coupled and decoupled paradigms.
7.1 Model Dynamics and Probabilistic Factorization
Let $R=(r_{1},r_{2},\dots,r_{T})$ be the sequence of latent reasoning states and $X=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$ be the sequence of generated tokens.
Classical Transformer Model.
A standard Transformer model $\mathcal{M}_{t}$, implicitly defines a joint probability distribution that is factorized sequentially. The generation of the state and token at step $t$ depends on both the state and the sampled token from step t1t-1:
$P(R,X)=\prod_{t=1}^{T}P(r_{t},\hat{x}_{t}|r_{t-1},\hat{x}_{t-1})$
In this formulation, the distribution for the next reasoning state $r_{t}$ is directly conditioned on the previously sampled token $\hat{x}_{t-1}$. Consequently, a sampling error at step t1t-1 (i.e., $\hat{x}_{t-1}\neq x^{*}_{t-1}$) introduces a persistent error into the reasoning state trajectory. This error corrupts the foundation for all subsequent reasoning and generation steps, leading to compounding error.
COCONUT Model.
The COCONUT model $\mathcal{M}_{c}$ is a coupled model with latent generation ability. It generates latent reasoning tokens before producing final tokens (coconut). Limited by the coupled architecture, latent vectors and tokens are arranged in the same sequence $Z=(z_{1},z_{2},\dots,z_{N})$, where each element $z_{t}$ can be either a continuous latent vector $r_{t}$ or a discrete token $\hat{x}_{t}$. Despite the different output types, the model follows the same autoregressive principle: every new element is conditioned on all prior elements.
$P(Z)=\prod_{t=1}^{N}P(z_{t}|z_{<t})$
Let’s consider a generation length of $T_{1}$ for latent reasoning, followed by $T_{2}$ tokens. The process unfolds as follows:
  • For the latent reasoning steps ($t=1,\dots,T_{1}$), the model generates $z_{t}=r_{t}$, conditioning on the previous latent vectors $z_{<t}=(r_{1},\dots,r_{t-1})$.
  • For the token generation steps ($t=T_{1}+1,\dots,T_{1}+T_{2}$), the model generates $z_{t}=\hat{x}_{t-T_{1}}$, conditioning on the full history of all preceding latent vectors and tokens, $z_{<t}=(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{t-T_{1}-1})$.
The unified sequence is the model’s critical limitation. Suppose the model has finished generating its reasoning chain $(r_{1},\dots,r_{T_{1}})$ and the first error appeared at the $n^{th}$ token, $\hat{x}_{n}\neq x^{*}_{n}$. For the very next step, $t=T_{1}+n+1$, the model make prediction based on the history $(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{n})$. The erroneous token $\hat{x}_{n}$ is now an immutable part of the model’s context, corrupting every subsequent decision. The error propagation is direct and unavoidable because reasoning and generation are inextricably linked in the same autoregressive sequence.
Decoupled Model $\mathcal{M}_{d}$:
In contrast, our JEPA-Reasoner architecture imposes a structural constraint on the generative process, yielding a more robust factorization where the reasoning chain is generated independently of the token sampling:
$P(R,X)$ $=P(R)\cdot P(X|R)$
$=\left(\prod_{t=1}^{T}P(r_{t}|r_{t-1})\right)\cdot\left(\prod_{t=1}^{T}P(\hat{x}_{t}|R,\hat{x}_{1:t-1})\right)$
This factorization reveals two key theoretical advantages.
  1. 1.
    Error Containment: The reasoning trajectory’s probability, $P(R)$, is independent of the token generation process $P(X|R)$. An error in sampling a token $\hat{x}_{t-1}$ has no mathematical pathway to influence the reasoning trajectory $R$. The high-level plan remains intact and stable. Furthermore, the normalization of reasoning vectors $r_{t}$ to the unit hypersphere ensures this trajectory is inherently bounded, preventing error amplification within the reasoning dynamics itself.
  2. 2.
    Mechanism for Recovery: At every step $t$, the token generator $P(\hat{x}_{t}|\cdot)$ is conditioned on the entire, lossless reasoning chain $R$. This provides a strong, stable signal that allows the Talker to potentially recover from a local token error in its own history ($\hat{x}_{1:t-1}$), an effect empirically validated in our ablation study (Appendix C).
This inherent error containment and recovery mechanism explains the superior robustness observed in our CFG experiments (Section 5.2).
@@ 8 Summary
We introduce JEPA-Reasoner, a novel architecture that decouples latent space reasoning from token generation. Our approach enables continuous latent reasoning guidance while mitigating step-by-step error propagation. Efficient parallel training was also made possible without sacrificing latent reasoning performance compared with COCONUT. Our experiments on synthetic tasks suggest that by decoupling the high-level latent space reasoning process from low-level token generation, JEPA-Reasoner produced promising potential for multi-threaded reasoning and exhibited enhanced robustness to input noise and error accumulation when generating structured sequences.
@@ B.1 CFG Rules and Sample
CFG can hierarchically produce highly structured expressions by replacing non-terminal symbols at each level with next-level symbols following a production rule, as shown in Figure 4. A sequence of terminal symbols is considered to be valid if it can be transformed back to the root symbol with dynamic programming and the given production rule. The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic. We designed our own CFG following the method used by (allen-physics). The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences.
@@
Previous research (allen-physics) shows that Transformer blocks can encode the structure of CFG rules within parameters. We assume that a robust model should be able to recognize the high-level structure of the input sequence, thus ignoring faulty tokens in the input. Since each high-level element in our CFG sequence produces 3 to 4 tokens, the model should be able to maintain relatively stable performance across at least 4 generation steps.
@@ C.1 Experiment Setup
Using the training method mentioned in Section 4, we use the JEPA-Reasoner model initialized with Transformer blocks trained on C4 and Wikitext (wikitext-dataset) dataset in this experiment setup to produce a human-readable result. We conducted controlled experiments using two sample inputs from the training dataset to evaluate the dependency of the Talker model on the Reasoner’s output:
@@ Appendix D K Value in Scaled Cosine Distance Loss
Appendix E Visualization of Mixed Latent Vectors
To visualize that JEPA-Reasoner can produce mixed latent vectors, we gathered the embedding vectors and model predictions from our tree-search experiments in this section. We extracted embedding vectors from distinct tree leaves alongside the model’s output latent representations after one forward pass. Principal Component Analysis (PCA) was applied to the collected embeddings and model predictions, and the visualization focuses on the first two principal components (PC1 and PC2).
As demonstrated in the Figure 6, predicted latent vectors (blue points in the figure) form a continuous cloud within the space spanned by discrete vocabulary embeddings (red diamond shapes). This distribution supports the experiment results that they are the linear combinations of vocabulary embeddings. Also, the predicted vectors do not converge to singular vocabulary points, providing empirical evidence for the hypothesis that JEPA-Reasoner is capable of maintaining information from multiple possible choices rather than committing to a single answer.

### v3

@@ 5.2.3 Model Configurations and Training Methods
We define three model variants: the vanilla Transformer baseline ($T$), the coupled COCONUT model ($C$), and our decoupled JEPA-Reasoner ($R$). Table 3 shows detailed model configurations used in this section:
@@
We first pretrain a Transformer on CFG data using cross-entropy loss. The checkpoint with the highest accuracy is selected as the initialization for all subsequent models  [^1] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count.. We then post-train $T$ as a baseline. $C$ is trained to predict a sequence of hidden states followed by tokens. $R$ is trained following the methodology in Section 4. All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, batch size 128, and context length 1024) and all models are trained until the loss stabilizes. Then the best-performing checkpoint is selected as the representative for each architecture.
@@ 5.2.4 Results and Conclusions
The results of the above experiments confirm the robustness advantage of the decoupled architecture. In the token-level error experiment, JEPA-Reasoner demonstrated less performance degradation than the baseline ($T$) when facing input noise during multi-step CFG completion tasks (Figure 3). JEPA-Reasoner also exhibits higher performance across different magnitudes of Gaussian noise in the latent space error experiment (Table 4), providing more empirical evidence for its robustness advantage.
@@
Table 4: Exact match accuracy of JEPA-Reasoner $R$ and COCONUT $C$ under different latent noise levels.
@@
These results demonstrate that JEPA-Reasoner has the potential to address the limitations of existing paradigms discussed in Section 2. By operating in a normalized latent space and offloading token generation to the Talker module, subsequent reasoning outputs do not condition on previous decisions, thereby mitigating error accumulation in the autoregressive process and enabling robust sequential generation under noisy conditions.
@@ 6 Real World Evaluation
We evaluate the 694M JEPA-Reasoner and 198M Mono-Talker (0.9B combined) on natural language tasks following the training methods in Section 4. Our analysis focuses on two key comparisons:
@@
    Comparison with Token-Level Model: The performance difference between the fine-tuned Transformer model (token level prediction) and JEPA-Reasoner paired with its Mono-Talker model (latent level prediction).
@@
    Comparison with Other Reasoning Paradigms: The performance of JEPA-Reasoner relative to other latent and non-latent reasoning models of similar scale.
6.1 Model Training
We first pretrain a standard 0.9B Transformer model on C4 and Wikitext (Merity et al., 2016) for 300k steps, then fine-tune it for 42k steps using general math QA pairs. The best-performing checkpoint is selected to initialize and further fine-tune a Transformer model to obtain the baseline for comparison. We also trained a JEPA-Reasoner + Talker pair following the methodology in Section 4 with the same data and hyperparameters used to obtain the Transformer baseline. The best-performing checkpoints of each architecture that emerged during each training phase are selected as their representative.
6.2 Performance Gain from Decoupled Architecture
We evaluated both the fine-tuned Transformer model and JEPA-Reasoner on the GSM8K (Cobbe et al., 2021) benchmark. Table 5 shows the performance under both 5-shot and 8-shot settings.
Table 5: GSM8K performance of Transformer baseline and JEPA-Reasoner
@@ 6.2 Performance Gain from Decoupled Architecture
Accuracy Gain.
JEPA-Reasoner substantially outperforms the fine-tuned Transformer, achieving performance gains of 79.2% (5-shot) and 149.5% (8-shot) on the GSM8K benchmark compared with the baseline. These results confirm that adapting to a decoupled, latent-space reasoning paradigm significantly improves the model’s ability to utilize internally learned knowledge.
Notably, both the Transformer baseline and JEPA-Reasoner are trained with identical data and hyperparameters to ensure fair comparison. Furthermore, the scaled cosine similarity loss used in SST prioritizes smooth latent transitions over logical correctness. Together, these factors suggest that the performance gain stems from the advantage of the decoupled latent reasoning architecture rather than new knowledge obtained during SST.
@@
Crucially, the JEPA-Reasoner exhibits superior scaling behavior. While the Base Transformer’s performance stagnates between 5-shot and 8-shot settings (improving only by 0.1%), the JEPA-Reasoner utilizes the additional examples to achieve a 131.7% improvement in performance. This scaling behavior indicates that the decoupled latent space reasoning architecture effectively overcomes the reasoning plateau, which was widely observed in small-scale token-based models, allowing for more robust logical deduction that is less constrained by surface-level token statistics.
6.3 Performance Comparison with Other Reasoning Paradigms
We evaluate JEPA-Reasoner against established reasoning paradigms like Chain-of-Thought (CoT) and Recurrent Depth. Table 6 shows the GSM8K benchmark performance of our JEPA-Reasoner model (paired with Talker) relative to other models of comparable size or performance [^2] Scores for Llama 3.2-1B (Meta AI, 2024) and Huginn-0125 (Geiping et al., 2025) are cited from their official reports. As official 8-shot scores are unavailable for Qwen 3 0.6B, we evaluated it using lm-evaluation-harness (Gao et al., 2024), employing the same configuration used for JEPA-Reasoner..
Comparison of JEPA-Reasoner against other reasoning paradigms on the GSM8K benchmark. Note that the COCONUT model is excluded as public 8-shot results are unavailable, and prior work indicates it trades performance for efficiency relative to CoT (Hao et al., 2024).
Table 6: Comparison of JEPA-Reasoner against other reasoning paradigms on the GSM8K benchmark. Note that the COCONUT model is excluded as neither its checkpoints nor its 8-shot GSM8K benchmark result are publicly available, and prior work indicates it trades performance for efficiency relative to CoT (Hao et al., 2024).
@@ 6.3 Performance Comparison with Other Reasoning Paradigms
Decoupled Latent Reasoning JEPA-Reasoner (Ours) 0.9B 51.9
@@
As illustrated in Table 6, JEPA-Reasoner demonstrates superior performance on the GSM8K benchmark compared to both CoT models and other latent reasoning architectures. With a parameter count of only 0.9B, our model achieves an 8-shot score of 51.9%, outperforming the CoT baseline of the same size (Llama 3.2-1B) by 7.5 percentage points. While the Recurrent Depth model (Huginn-0125) offers a strong baseline at 42.1%, it requires nearly four times the parameter count to achieve results that are still 9.8 percentage points lower than JEPA-Reasoner. Consequently, these results serve as strong empirical evidence that JEPA-Reasoner’s decoupled latent reasoning architecture is capable of handling complex natural language reasoning tasks both effectively and efficiently.
@@ 7 Theoretical Advantage of Decoupled Architecture
the robustness advantage of JEPA-Reasoner is built on two complementary mechanisms: the decoupled architecture design and the normalized property of latent space. We can formalize this advantage by analyzing the probabilistic assumptions and information flow within the coupled and decoupled paradigms. Denote $R=(r_{1},r_{2},\dots,r_{T})$ as the sequence of latent reasoning states and $X=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$ as the sequence of generated tokens.
7.1 Decoupled Architecture Stops the Propagation of Error
Assume that the joint probability can be factored into $P(R,X)=P(R)\cdot P(X|R)$. The autoregressive update rules for the reasoning state $r_{t}$ and token $\hat{x}_{t}$ at step $t$ are defined as:
$\begin{cases}r_{t}=f_{\theta}(r_{t-1})&\text{Reasoner Update}\\ \hat{x}_{t}\sim g_{\phi}(r_{t},\hat{x}_{1:t-1})&\text{Talker Sampling}\end{cases}$
where $f_{\theta}$ is the deterministic predictor function of the Reasoner and $g_{\phi}$ is the probabilistic decoding function of the Talker. We define the sensitivity of the future reasoning state $r_{t+1}$ with respect to the current generated token $\hat{x}_{t}$ using partial derivatives as follows. Noticing that by the definition of decoupled architecture, $r_{t}$ depends solely on $r_{t-1}$. Considering that the generated token $\hat{x}_{t}$ does not appear in the input argument of $f_{\theta}$. Which means that $\frac{\partial r_{t}}{\partial\hat{x}_{\tau}}=0\quad\forall\tau\geq 1$. Using the chain rule, we obtain:
$\frac{\partial r_{t+1}}{\partial\hat{x}_{t}}=\frac{\partial f_{\theta}(r_{t})}{\partial r_{t}}\cdot\underbrace{\frac{\partial r_{t}}{\partial\hat{x}_{t}}}_{0}=\mathbf{0}$
As a conclusion, there is no mathematical pathway for token-space errors produced during sampling to propagate into the latent reasoning trajectory, which enables better robustness during generation
7.2 Bounded Error via Normalization and Angular Alignment
In this section, we demonstrate that unlike standard autoregressive models where errors can compound in both magnitude and direction, JEPA-Reasoner’s error bound is explicitly enforced by the L2 normalization layer and implicitly minimized by the angular training objective. This dual constraint prevents the reasoning trajectory produced from diverging unbounded, providing the theoretical basis for the superior robustness observed in our experiments.
Explicitly Constrained by Architecture
By constraining all latent states to the unit hypersphere $\mathcal{S}^{d-1}$, the divergence between an arbitrary trajectory produced by JEPA-Reasoner during runtime (denoted by $R$) and an optimal reference trajectory (denoted by $R^{*}$) is strictly bounded. Since the predictor output is passed through an L2 normalization layer before looping back, every reasoning state $r_{t}$ satisfies $\|r_{t}\|_{2}=1$. Consequently, the Euclidean distance between the predicted state and the optimal state is physically constrained by the diameter of the hypersphere:
$\|r_{t}-r_{t}^{\prime}\|_{2}\leq\text{diam}(S^{d-1})=2$
Implicitly Minimized by Training Objective
Furthermore, the scaled cosine distance loss $\mathcal{L}_{t}\propto 1-\cos(r_{t},r^{*}_{t})=\frac{1}{2}\|r_{t}-r^{*}_{t}\|_{2}^{2}$ used in our Self-Supervised Training (SST) phase encourages the model to be aligned with a continuous trajectory on a bounded hypersphere instead of making leaps between discrete tokens like a traditional Transformer, implicitly reducing the risk of producing arbitrary latent vectors that diverges from the optimal reasoning path.
@@ 8 Summary
We introduce JEPA-Reasoner, a novel architecture that decouples latent space reasoning from token generation. Our approach enables continuous latent reasoning guidance while mitigating step-by-step error propagation. Efficient parallel training was also made possible without sacrificing latent reasoning performance compared with COCONUT. Our experiments suggest that by decoupling the high-level latent space reasoning process from low-level token generation, JEPA-Reasoner produced promising potential for multi-threaded reasoning and exhibited enhanced robustness to input noise and error accumulation when generating structured sequences.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 CFG Rules and Sample
CFG can hierarchically produce highly structured expressions by replacing non-terminal symbols at each level with next-level symbols following a production rule, as shown in Figure 5. A sequence of terminal symbols is considered to be valid if it can be transformed back to the root symbol with dynamic programming and the given production rule. The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic. We designed our own CFG following the method used by (Allen-Zhu and Li, 2023). The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences.
@@
Previous research (Allen-Zhu and Li, 2023) shows that Transformer blocks can encode the structure of CFG rules within parameters. We assume that a robust model should be able to recognize the high-level structure of the input sequence, thus ignoring faulty tokens in the input. Since each high-level element in our CFG sequence produces 3 to 4 tokens, the model should be able to maintain relatively stable performance across at least 4 generation steps.
@@ C.1 Experiment Setup
Using the training method mentioned in Section 4, we use the JEPA-Reasoner model initialized with Transformer blocks trained on C4 and Wikitext (Merity et al., 2016) dataset in this experiment setup to produce a human-readable result. We conducted controlled experiments using two sample inputs from the training dataset to evaluate the dependency of the Talker model on the Reasoner’s output: